# Language-Codec: Reducing the Gaps Between Discrete Codec Representation and Speech Language Models

### v3

@@ 4.3. Generalization Evaluation
We validated the generalization performance of the codec model on a total of 13,100 audio samples from the LJSpeech dataset. The audio in the LJSpeech dataset has a sampling rate of 22,050 Hz, which we resampled to 24,000 Hz during the input stage of the inference process. Since the codec model was trained on tens of thousands of hours of speech data, it possesses a stronger generalization capability. From Table 2, we observed the following findings: 1) Most codec models demonstrated impressive generalization performance, with SpeechTokenizer exhibiting slightly lower generalization performance, likely due to training data limitations. 2) In tests involving different sampling rates and out-of-domain samples, Language-Codec outperformed the current SOTA baseline models significantly across various objective metrics such as PESQ, STOI, V/UV F1, and SPK, for both the four-channel and eight-channel configurations. For the objective metric UTMOS, Language-Codec maintained comparable performance to the current SOTA models. In general, due to its data-driven nature, Language-Codec demonstrates excellent generalization performance.
4.4. Zero-shot TTS Evaluation
Regarding the inference phase, to ensure fair comparisons, we followed the experimental protocols outlined in VALL-E (Wang et al., 2023) and employed the LibriSpeech Test-Clean dataset (Panayotov et al., 2015), ensuring no overlap with our training data. We specifically utilized samples from the LibriSpeech Test-Clean dataset with durations ranging from 4 to 10 seconds, resulting in a subset of 2.2 hours. Following VALL-E, we use the whole transcription and the first 3 seconds of the utterance as the phoneme and acoustic prompts respectively, and ask the model to generate the continuations.
Table 3. Evaluation of zero-shot TTS models with different codecs on the LibriSpeech Test-Clean corpus.
Zero-shot TTS Model WER $\downarrow$ SPK $\uparrow$ MOS-Q $\uparrow$ MOS-P $\uparrow$ MOS-S $\uparrow$ VALL-E w/ Encodec 4.3 0.6115 3.73$\pm$0.09 3.76$\pm$0.12 3.74$\pm$0.11 VALL-E w/ Encodec+Vocos 4.3 0.6198 3.83$\pm$0.07 3.81$\pm$0.13 3.82$\pm$0.06 VALL-E w/ Language-Codec 4.1 0.6806 3.93$\pm$0.11 3.84$\pm$0.09 3.96$\pm$0.08 MobileSpeech w/ Encodec 3.2 0.6776 3.91$\pm$0.10 3.99$\pm$0.12 3.98$\pm$0.11 Mobilespeech w/ Encodec+Vocos 3.1 0.6883 4.05$\pm$0.08 4.03$\pm$0.12 4.04$\pm$0.09 MobileSpeech w/ Language-Codec 3.0 0.7544 4.12$\pm$0.12 4.03$\pm$0.08 4.11$\pm$0.07
Due to the prevalent utilization of Encodec models in downstream speech language models to extract features as intermediate acoustic tokens, we have employed it as the primary baseline model. In the case of the Vocos version, we continue to train downstream models using representations extracted by the Encodec model, but during the inference decoding phase, we employ the Vocos decoder. Detailed explanations will be provided in the following section on ablation experiments. Furthermore, for the Language-Codec version, we exclusively train using codec representations extracted by the Language-Codec model and utilize the Language-Codec model for inference.
[…]
Moreover, during the training process of our downstream zero-shot Text-to-Speech model, we find that when the downstream model predicts codecs generated by the Language-Codec model, the accuracy of codec prediction decreases when the number of channels exceeds four. Although this does not have a significant impact on the performance of the downstream model, future endeavors could explore the use of smaller or variable codebooks to further enhance the results.
4.5. Ablation experiment
In this section, we conducted a detailed analysis of the Vocos module and MCRVQ module. Furthermore, regarding the significant improvement of Language-Codec over the current SOTA baselines, this outcome is attributed to several factors, including the corresponding training data volume and data quality, the selection of hyperparameters (such as model depth, loss coefficients, optimizer configurations, etc.), and the joint optimization of multiple adversarial losses. We have made available the best set of hyperparameter configurations as an open-source resource.
Vocos Modules
We conducted ablation tests on the ConvNeXt Blocks and Fourier Transform structure of the decoder to evaluate its impact on the codecs reconstruction. Specifically, we employed the pre-trained encoder module, the quantizer module from Encodec and the pre-trained decoder module from Vocos for inference. In Tables 1, we refer to this model as Vocos. By comparing the Encodec model and the Vocos model on the LibriTTS Test-Clean and Test-Other datasets, we observed that replacing the decoder with the Vocos structure significantly improves the audio quality in the four-channel setting. However, in the eight-channel setting, the F1 and SPK scores of the Vocos model were slightly lower than those of the Encodec model, although the UTMOS, PESQ, and STOI scores were higher. Subjectively, the Vocos structure effectively mitigates the artifacts introduced by the Encodec model. Furthermore, when comparing the Vocos model with the Language-Codec model, we found that the Language-Codec model outperforms the Vocos model significantly. This observation suggests that the MCRVQ mechanism, which normalizes the information in the quantizer, further enhances the audio reconstruction quality.
MCRVQ
We also validated the role of the Masked Channel Residual Vector Quantization (MCRVQ) module in the language-codec model. Considering that the design purpose of the MCRVQ mechanism is to reduce the difficulty of text generation in downstream tasks, we conducted ablation experiments on the zero-shot TTS model downstream. Specifically, we first replaced the MCRVQ module in the codec model with the RVQ module in the encodec model while keeping the same training steps and other configurations. We refer to this experiment setup as Language-Codec w/o MCRVQ. We used Language-Codec w/o MCRVQ to extract the corresponding discrete codec features and retrained the downstream VALL-E and MobileSpeech models. The experimental results, as shown in Table 4, revealed that there was no significant difference between Language-Codec w/o MCRVQ and Language-Codec in terms of the robustness metric WER. However, in terms of the objective metric of speaker similarity, omitting the MCRVQ module resulted in a decrease of 0.04 and 0.05 similarity in VALL-E and MobileSpeech, respectively, indicating that the MCRVQ module indeed enhances the codec generation capability of the downstream speech synthesis model by weakening the difficulty of text generation for codec. In addition, we also conducted corresponding subjective CMOS tests, from Table 5, it can be observed that in the autoregressive discrete codec modeling experiments of the VALL-E model, the CMOS values of the synthesized audio decreased by 0.13 when the MCRVQ module was omitted compared to the original Language-Codec model. Similarly, in the parallel discrete codec modeling experiments of the MobileSpeech model, the CMOS values of the synthesized audio decreased by 0.17 when the MCRVQ module was omitted compared to the original Language-Codec model, which further indicated that the codec generated by the Language-Codec w/o MCRVQ model had lower subjective audio quality and audio similarity than the codec generated by the Language-Codec model.
Table 4. The ablation experiments of the MCRVQ module, we assessed the performance of WER and SPK metrics.
Model Codec Model WER $\downarrow$ SPK $\uparrow$ VALL-E Language-Codec w/o MCRVQ 4.2 0.6403 VALL-E Language-Codec 4.1 0.6806 MobileSpeech Language-Codec w/o MCRVQ 3.1 0.7085 MobileSpeech Language-Codec 3.0 0.7544
Table 5. The ablation experiments of the MCRVQ module, we assessed the performance of subjective metric CMOS.
Model Codec Model CMOS $\uparrow$
VALL-E Language-Codec w/o MCRVQ -0.13
VALL-E Language-Codec 0.00
MobileSpeech Language-Codec w/o MCRVQ -0.17
MobileSpeech Language-Codec 0.00
5. Conclusion and Limitations
In this article, we propose Language-Codec, a discrete acoustic codec model that enhances adaptation to downstream speech language models. Through larger training data, improved model architecture, better hyperparameters, and a unique Masked Channel residual vector quantization mechanism, we achieve excellent audio reconstruction quality with just four layers of codecs. The Language-Codec model demonstrates effective audio restoration performance in both clean audio and noisy environments. Furthermore, we validate the generalization capability of Language-Codec in unseen domains and its generation ability in downstream zero-shot TTS models, yielding satisfactory results. We envision Language-Codec as a state-of-the-art foundational codec model for future research in the field of speech generation.

### v4

@@ Appendix E More Reconstruction Evaluation
We validated the generalization performance of the codec model on a total of 13,100 audio samples from the LJSpeech dataset. The audio in the LJSpeech dataset has a sampling rate of 22,050 Hz, which we resampled to 24,000 Hz during the input stage of the inference process. Since the codec model was trained on tens of thousands of hours of speech data, it possesses a stronger generalization capability. From Table 7, we observed the following findings: 1) Most codec models demonstrated impressive generalization performance, with SpeechTokenizer exhibiting slightly lower generalization performance, likely due to training data limitations. 2) In tests involving different sampling rates and out-of-domain samples, Language-Codec outperformed the current SOTA baseline models significantly across various objective metrics such as UTMOS, PESQ, STOI, V/UV F1, and SPK, for both the four-channel and eight-channel configurations. In general, due to its data-driven nature, Language-Codec demonstrates excellent generalization performance.
Appendix F Ablation experiments about discriminator
We evaluated the impact of the ablation experiments on the multi-scale discriminator on the LibriTTS testclean dataset. The experimental results are presented in Table 8. We observed that the multi-scale discriminator enhances the reconstruction performance of the Codec model to a certain extent. Specifically, the multi-scale discriminator shows significant improvements in the PESQ metric.
Appendix G Ablation experiments about decoder
We conducted detailed ablation experiments on the Decoder module within the Language-Codec framework. Specifically, we initially replaced the Decoder with the upsampling module from the Encodec model. This framework was denoted as ’Language-Codec w/o decoder’. Additionally, we examined the impact of the attention module on Codec reconstruction fidelity. This framework was denoted as ’Language-Codec w/o attention block’. The experimental results, as shown in Table 9, highlight the significant influence of the upsampling structure dependent on Fourier transforms and the attention module on Codec reconstruction efficacy.
Appendix H Experiments on downstream generative models with four quantizers
Due to time constraints, we evaluated the generation performance of a four-layer quantizer on the MobileSpeech model. The experimental results, presented in Table 10, demonstrate that even with just four codebooks, high-quality audio can be generated.
Appendix I Visualize modules in Decoder
As shown in Figure 2(a), the Attention Block is composed of stacked Conv1d ResBlocks and Self-Attention modules.
Specifically, as shown in Figure 2(b), The Conv1d ResBlock processes the features by first applying normalization and sigmoid activation, followed by a Conv1d layer with the kernel size of 3, stride of 1, and padding of 1. After another round of normalization and sigmoid activation, the features undergo dropout before being passed through another Conv1d layer with the same configuration. Finally, residual connections are applied to produce the output.
The self-attentive module, as shown in Figure 2(c), utilizes three Conv1d layers with kernel size of 1, stride of 1, and padding of 0 to extract Q, K, and V, respectively. After that execute the attention mechanism and finally go through the Conv1d layer output with kernal size of 1, Stride of 1 and padding of 0.
As shown in Figure 2(d), ConvNeXt consists of the stack of N ConvNeXtBlocks, each ConvNeXtBlock processes the input features using the Conv1d layer with Kernal size of 7, Stride of 3, and padding of 1. After Normalize and Linear layer, the output features are processed by GLUE activation and the Linear layer, finally using the residual connection for the input of the next ConvNeXtBlock.